OpenStudy (anonymous):

what is the least common denominator x, x+1, 3X HELP PLEASE

OpenStudy (anonymous):

the product of these three

OpenStudy (itiaax):

The least common denominator is x

OpenStudy (anonymous):

how so?

OpenStudy (itiaax):

Okay..let's say x represented 5? x=5 x+1=6 3x=15

OpenStudy (amoodarya):

lcm depends on value of x but lcm [x,x+1,3x]=3x(x+1)

OpenStudy (mathstudent55):

LCD is obtained by looking at the factors of each quantity. x = x x + 1 = x + 1 3x = 3 * x You need one factor of each. Use x, x + 1 and 3, so LCD is 3x(x + 1)

OpenStudy (amoodarya):